OpenStudy (anonymous):

can someone PLEASSEE help explain this problem to me? It's multiplying polynomials (x^4+4) (x^2-4)

OpenStudy (mertsj):

Use FOIL

OpenStudy (anonymous):

(x^4 * X^2) - (4 *4) +(4x^2)-(4x^4)

OpenStudy (anonymous):

= X^6 -4X^4 +4X^2-16

OpenStudy (anonymous):

I haven't learned about foil yet but thank you. and thanks @adel_demo . do you think there's anyway you could explain to me the process of how you did that

OpenStudy (anonymous):

sure one second

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

multiply the outer elements and add them to the inner elements then add them to the multiplication of the first element and the first element then add this to the multiplication of the second element and the second element

OpenStudy (anonymous):

and of course dont forget the sign rule + * + = + - * - = + _ * + = -

OpenStudy (anonymous):

thank you so much!!!!!! i have one more question - (x^2+3x+1) (x^2+x+2) i got a bunch of different answers but this is the format for me to type in for the answer- "_ _ + _ _ _ + _ _ _ + _ _ _ _"

OpenStudy (anonymous):

@adel_demo

OpenStudy (anonymous):

now we have a diffrant type and we need more than 4

OpenStudy (anonymous):

take the first element x^2 and multiply it to the (x^2+x+2)

OpenStudy (anonymous):

and then take the second element 3x and multiply it to (x^2+x+2)

OpenStudy (anonymous):

then take the third element 1 and multiplay it to (x^2+x+2)

OpenStudy (anonymous):

then u would have 9 new elements add them to gether u should get a 6 elements only at the end

OpenStudy (anonymous):

okay so i did the first element to start and got this- 2x^4+1x^2_2^2 did i do it right?

OpenStudy (anonymous):

i didn't mean to start that with a - sign

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

see X^2 * X^2 = X^4 not 2X^4

OpenStudy (anonymous):

try again :)

OpenStudy (anonymous):

ohh! i forgot to add the powers! so is it x^4+x^2+2x^2 for the first term?

OpenStudy (anonymous):

for the 2nd term is it- 3x^2 + 3x+6

OpenStudy (anonymous):

the middle one is powe 3 not 2

OpenStudy (anonymous):

it is 3X not three this is ur mistake

OpenStudy (anonymous):

i mean the middle element of the first group

OpenStudy (anonymous):

ok thanks i understand. this is what i have now 1st term: x^4+x^3+2x^2 2nd term: 3x^3+3x^2+6x 3rd term: 1x^2+1x+2

OpenStudy (anonymous):

wow this is wonderful ^_^ now add the ones with the same power together :)

OpenStudy (anonymous):

this is what i got but i wasn't sure if i was supposed to add the actual powers together. x^4+4x^3+6x^2+7x+2

OpenStudy (anonymous):

this is soo right :)

OpenStudy (anonymous):

omg!! thanks soo much! :)

OpenStudy (anonymous):

ur welcome :)